OpenStudy (anonymous):

solve for x please.....3x-8=-x+4

OpenStudy (anonymous):

3x+x-8=4 4x-8=4 4x=12 x=3

OpenStudy (anonymous):

thankk youuu

OpenStudy (anonymous):

no problem ;)

OpenStudy (anonymous):

3x+x=4+8 4x=12 divide both sides by 4 x=3 :)